OpenStudy (anonymous):

Double checking work here. Solve 15x^4-14x^2+3=0 My answers are \[1/\sqrt{3}, - 1/\sqrt{3}\] \[\sqrt{3/5}, -\sqrt{3/5}\]

OpenStudy (apoorvk):

the last two are right the first two will be sqrt3 and -sqrt3

OpenStudy (anonymous):

so i towuld be \[\sqrt{3}, -\sqrt{3}\] for the first two

OpenStudy (apoorvk):

yes exactly... re do it... at the factor (x^2-3)=0, i guess u made a mistake.

OpenStudy (anonymous):

yea i think i messed up the fraction, inverted them probably on accident

OpenStudy (anonymous):

no i actually double checked and the first part ± \[\sqrt{1/3}\]

OpenStudy (anonymous):

±